Hello my name is Lauren Blue and I am trying to study for my final on Wednesday June 13th, 2012 and I am having trouble with the Multiplying and Dividing Radical Expressions. THe question I am stuck on is (4 multiplied by the square root of x) multiplied by (-2 multiplied by the square root of x).
Answer by richwmiller(9135)

4*sqrt(x)*-2*sqrt(x)
think of sqrt(x) as y
4*y*(-2)*y=-8y^2
since y^2=(sqrt(x))^2=x equals
-8y^2=-8*x